# Double Robustness of Local Projections

### v1

@@ Appendix B Proofs of propositions
$\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{h,\ell}^{\prime}\Xi_{h,\ell})=\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h}).$
Indeed, by multiplying out terms and using $\Sigma=HDH^{\prime}$, we obtain
$\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{h,\ell}^{\prime}\Xi_{h,\ell})$ $=\sum_{\ell=1}^{\infty}\operatorname*{trace}(A^{\ell-1}\Sigma(A^{\prime})^{\ell-1}S^{-1}\Psi_{h}\Sigma\Psi_{h}^{\prime}S^{-1})$
$\quad-2\sum_{\ell=1}^{h}\operatorname*{trace}(A^{\ell-1}H_{\bullet,j^{*}}e_{i^{*},n}^{\prime}A^{h-\ell}\Sigma\Psi_{h}^{\prime}S^{-1})$
$\quad+\sum_{\ell=1}^{h}\sigma_{j^{*}}^{-2}e_{i^{*},n}^{\prime}A^{h-\ell}\Sigma(A^{\prime})^{h-\ell}e_{i^{*},n}.$
The first term on the right-hand side above equals
$\operatorname*{trace}\bigg{\{}\underbrace{\left(\sum_{\ell=1}^{\infty}A^{\ell-1}\Sigma(A^{\prime})^{\ell-1}\right)}_{=\operatorname*{Var}(\tilde{y}_{t})=S}S^{-1}\Psi_{h}\Sigma\Psi_{h}^{\prime}S^{-1}\bigg{\}}=\operatorname*{trace}\{\Psi_{h}\Sigma\Psi_{h}^{\prime}S^{-1}\}.$
The second term on the right-hand side two displays ago equals
$-2\operatorname*{trace}\bigg{\{}\underbrace{\left(\sum_{\ell=1}^{h}A^{\ell-1}H_{\bullet,j^{*}}e_{i^{*},n}^{\prime}A^{h-\ell}\right)}_{=\sum_{\ell=1}^{h}A^{h-\ell}H_{\bullet,j^{*}}e_{i^{*},n}^{\prime}A^{\ell-1}=\Psi_{h}}\Sigma\Psi_{h}^{\prime}S^{-1}\bigg{\}}=-2\operatorname*{trace}\{\Psi_{h}\Sigma\Psi_{h}^{\prime}S^{-1}\}.$
We have shown
$\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{h,\ell}^{\prime}\Xi_{h,\ell})=\frac{1}{\sigma_{j^{*}}^{2}}\sum_{\ell=1}^{h}e_{i^{*},n}^{\prime}A^{h-\ell}\Sigma(A^{\prime})^{h-\ell}e_{i^{*},n}-\operatorname*{trace}\{\Psi_{h}\Sigma\Psi_{h}^{\prime}S^{-1}\},$
which in turn equals $\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})$ (cf. 3.1), as required. ∎
Appendix C Proofs of corollaries
C.1 Proof of 3.1
By Proposition 3.1 and the that fact $\xi_{h,i^{*},t}\varepsilon_{j^{*},t}$ is serially uncorrelated, the asymptotic variance of LP equals $\sigma_{j^{*}}^{-4}E[\xi_{h,i^{*},t}^{2}\varepsilon_{j^{*},t}^{2}]=\sigma_{j^{*}}^{-2}E[\xi_{h,i^{*},t}^{2}]$, where $\xi_{h,i^{*},t}$ is the $i^{*}$-th element of the $n\times 1$ vector
with $\overline{H}_{j^{*}}\equiv(H_{\bullet,j^{*}+1},\dots,H_{\bullet,m})$ and $\overline{\varepsilon}_{j^{*},t}\equiv(\varepsilon_{j^{*}+1,t},\dots,\varepsilon_{m,t})^{\prime}$. A quick calculation verifies the expression stated in the corollary.
Using the limiting representation for the VAR estimator in Proposition 3.2, notice first that $T^{-1}\sum_{t=1}^{T}\xi_{0,t}\varepsilon_{j^{*},t}$ and $T^{-1}\sum_{t=1}^{T}\varepsilon_{t}\tilde{y}_{t-1}^{\prime}$ are asymptotically independent, since for any $i\in\{1,\dots,n\}$, $j\in\{1,\dots,m\}$, and $\ell\geq 0$,
$E[\xi_{0,i,t}\varepsilon_{j^{*},t}\tilde{y}_{t-\ell-1}\varepsilon_{j,t-\ell}]=E[\xi_{0,i,t}]E[\varepsilon_{j^{*},t}\varepsilon_{j,t-\ell}]E[\tilde{y}_{t-\ell-1}]=0.$
The asymptotic variance of the term
$\frac{1}{\sigma_{j^{*}}^{2}}e_{i^{*},n}^{\prime}A^{h}T^{-1}\sum_{t=1}^{T}\xi_{0,t}\varepsilon_{j^{*},t},$
which has serially uncorrelated summands, is given by
$\frac{1}{\sigma_{j^{*}}^{4}}e_{i^{*},n}^{\prime}A^{h}E[\xi_{0,t}\xi_{0,t}^{\prime}]E[\varepsilon_{j^{*},t}^{2}]A^{h}e_{i^{*},n}=\frac{1}{\sigma_{j^{*}}^{2}}e_{i^{*},n}^{\prime}A^{h}\overline{H}_{j^{*}}\overline{D}_{j^{*}}\overline{H}_{j^{*}}^{\prime}(A^{\prime})^{h}e_{i^{*},n}.$
The asymptotic variance of the term
$\operatorname*{trace}\left\{S^{-1}\Psi_{h}HT^{-1/2}\sum_{t=1}^{T}\varepsilon_{t}\tilde{y}_{t-1}^{\prime}\right\}=T^{-1/2}\sum_{t=1}^{T}\tilde{y}_{t-1}^{\prime}S^{-1}\Psi_{h}H\varepsilon_{t},$
which also has serially uncorrelated summands, equals
$\operatorname*{Var}\left(\tilde{y}_{t-1}^{\prime}S^{-1}\Psi_{h}H\varepsilon_{t}\right)$ $=E[\tilde{y}_{t-1}^{\prime}S^{-1}\Psi_{h}H\varepsilon_{t}\varepsilon_{t}^{\prime}H^{\prime}\Psi_{h}^{\prime}S^{-1}\tilde{y}_{t-1}]$
$=\operatorname*{trace}\left\{E[S^{-1}\Psi_{h}H\varepsilon_{t}\varepsilon_{t}^{\prime}H^{\prime}\Psi_{h}^{\prime}S^{-1}\tilde{y}_{t-1}\tilde{y}_{t-1}^{\prime}]\right\}$
$=\operatorname*{trace}\left\{S^{-1}\Psi_{h}HE[\varepsilon_{t}\varepsilon_{t}^{\prime}]H^{\prime}\Psi_{h}^{\prime}S^{-1}E[\tilde{y}_{t-1}\tilde{y}_{t-1}^{\prime}]\right\}$
$=\operatorname*{trace}\left\{S^{-1}\Psi_{h}HDH^{\prime}\Psi_{h}^{\prime}S^{-1}S\right\}$
$=\operatorname*{trace}\left\{\Psi_{h}\Sigma\Psi_{h}^{\prime}S^{-1}\right\}.$
The expression for the asymptotic variance of VAR follows.
Finally, for the formula of the asymptotic covariance, note that by Propositions 3.1 and 3.2,
$\hat{\beta}_{h}-\theta_{h,T}$ $=R_{1,T}+R_{3,T}+o_{p}(T^{-1/2}),$
$\hat{\delta}_{h}-\theta_{h,T}$ $=T^{-\zeta}\operatorname{aBias}(\hat{\delta}_{h})+R_{2,T}+R_{3,T}+o_{p}(T^{-\zeta}+T^{-1/2}),$
where
$R_{1,T}$ $\equiv\frac{1}{\sigma_{j^{*}}^{2}}T^{-1}\sum_{t=1}^{T}\left(\sum_{\ell=1}^{h}e_{i^{*},n}^{\prime}A^{h-\ell}H\varepsilon_{t+\ell}\right)\varepsilon_{j^{*},t},$
$R_{2,T}$ $\equiv\operatorname*{trace}\left\{S^{-1}\Psi_{h}HT^{-1}\sum_{t=1}^{T}\varepsilon_{t}\tilde{y}_{t-1}^{\prime}\right\},$
$R_{3,T}$ $\equiv\frac{1}{\sigma_{j^{*}}^{2}}e_{i^{*},n}^{\prime}A^{h}T^{-1}\sum_{t=1}^{T}\left(\overline{H}_{j^{*}}\overline{\varepsilon}_{j^{*},t}\right)\varepsilon_{j^{*},t}.$
The asymptotic covariances for the order $T^{-1/2}$ terms in the representation above is
$\operatorname{aCov}(R_{1,T}+R_{3,T},R_{2,T}+R_{3,T}).$
In the calculation of $\operatorname{aVar}(\hat{\delta}_{h})$, we have already shown that $\operatorname{aCov}(R_{3,T},R_{2,T})=0$. Analogous calculations also show that $\operatorname{aCov}(R_{3,T},R_{1,T})=0$. Thus,
$\operatorname{aCov}(R_{1,T}+R_{3,T},R_{2,T}+R_{3,T})=\operatorname{aCov}(R_{1,T},R_{2,T})+\operatorname{aVar}(R_{3,T}).$
We next show that $\operatorname{aCov}(R_{1,T},R_{2,T})=\operatorname{aVar}(R_{2,T})$. Note that
$\operatorname{aCov}(R_{1,T},R_{2,T})$ $=\frac{1}{\sigma_{j^{*}}^{2}}\sum_{s=-\infty}^{\infty}\operatorname*{Cov}\left(\varepsilon_{j^{*},t}\sum_{\ell=1}^{h}e_{i^{*},n}^{\prime}A^{h-\ell}H\varepsilon_{t+\ell},\operatorname*{trace}\left\{S^{-1}\Psi_{h}H\varepsilon_{t+s}\tilde{y}_{t+s-1}^{\prime}\right\}\right)$
$=\frac{1}{\sigma_{j^{*}}^{2}}\sum_{\ell=1}^{h}E\left[e_{i^{*},n}^{\prime}A^{h-\ell}H\varepsilon_{t+\ell}\varepsilon_{t+\ell}^{\prime}H^{\prime}\Psi_{h}^{\prime}S^{-1}(\varepsilon_{j^{*},t}\tilde{y}_{t+\ell-1})\right]$
$=\frac{1}{\sigma_{j^{*}}^{2}}\sum_{\ell=1}^{h}e_{i^{*},n}^{\prime}A^{h-\ell}HE\left[\varepsilon_{t+\ell}\varepsilon_{t+\ell}\right]^{\prime}H^{\prime}\Psi_{h}^{\prime}S^{-1}E\left[\varepsilon_{j^{*},t}\tilde{y}_{t+\ell-1}\right]$
$=\frac{1}{\sigma_{j^{*}}^{2}}\operatorname*{trace}\left\{\left(\sum_{\ell=1}^{h}E\left[\varepsilon_{j^{*},t}\tilde{y}_{t+\ell-1}\right]e_{i^{*},n}^{\prime}A^{h-\ell}\right)HDH^{\prime}\Psi_{h}^{\prime}S^{-1}\right\}$
$=\frac{1}{\sigma_{j^{*}}^{2}}\operatorname*{trace}\left\{\left(\sum_{\ell=1}^{h}A^{\ell-1}H_{\bullet,j^{*}}\sigma_{j^{*}}^{2}e_{i^{*},n}^{\prime}A^{h-\ell}\right)\Sigma\Psi_{h}^{\prime}S^{-1}\right\}$
@@ Appendix C Proofs of corollaries
$=\operatorname*{trace}\left\{\left(\sum_{\ell=1}^{h}A^{\ell-1}H_{\bullet,j^{*}}e_{i^{*},n}^{\prime}A^{h-\ell}\right)\Sigma\Psi_{h}^{\prime}S^{-1}\right\}.$
Finally, reversing the order of the terms in the sum above gives
$\sum_{\ell=1}^{h}A^{\ell-1}H_{\bullet,j^{*}}e_{i^{*},n}^{\prime}A^{h-\ell}=\sum_{\ell=1}^{h}A^{h-\ell}H_{\bullet,j^{*}}e_{i^{*},n}^{\prime}A^{\ell-1}=\Psi_{h}.$
Thus, we conclude that
@@
$\operatorname{aCov}(R_{1,T},R_{2,T})=\operatorname*{trace}\left\{\Psi_{h}\Sigma\Psi_{h}^{\prime}S^{-1}\right\}=\operatorname{aVar}(R_{2,T}).$
where the final equality follows from the derivation of $\operatorname{aVar}(\hat{\delta}_{h})$. This implies that
$\operatorname{aCov}(R_{1,T}+R_{3,T},R_{2,T}+R_{3,T})=\operatorname{aVar}(R_{2,T})+\operatorname{aVar}(R_{3,T})=\operatorname{aVar}(\hat{\delta}_{h}).\qed$
C.2 Proof of 3.2
The result is an immediate consequence of Propositions 3.1 and 3.2. For the sake of exposition, we present details below. Recall the definitions of the VAR and LP confidence intervals in 3.2, and let the asymptotic variances $\operatorname{aVar}(\hat{\beta}_{h})$ and $\operatorname{aVar}(\hat{\delta}_{h})$ correspond to the formulae in 3.1.
First, we show that $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-a$. Note that
$P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))$ $=P\left(-z_{1-a/2}\leq\frac{\sqrt{T}(\widehat{\beta}_{h}-\theta_{h,T})}{\sqrt{\operatorname{aVar}(\hat{\beta}_{h})}}\leq z_{1-a/2}\right).$
By Proposition 3.1 and 3.1,
[…]
We thus conclude that $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-a$.
Second, we characterize $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\delta}_{h}))$. Note that
$P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=P\left(-\frac{z_{1-a/2}}{\sqrt{T}}\leq\frac{(\widehat{\delta}_{h}-\theta_{h,T})}{\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}}\leq\frac{z_{1-a/2}}{\sqrt{T}}\right).$
By Proposition 3.2 and 3.1,
$\hat{\delta}_{h}-\theta_{h,T}=T^{-\zeta}\operatorname{aBias}(\hat{\delta}_{h})+\tilde{R}_{T}+o_{p}(T^{-\zeta}+T^{-1/2}),$
where
[…]
Consider the following three cases.
  1. 1.
    $\zeta\in(1/4,1/2)$: In this case
    $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))$ $=\lim_{T\to\infty}P\left(-\frac{z_{1-a/2}}{T^{1/2-\zeta}}\leq\frac{T^{\zeta}(\widehat{\delta}_{h}-\theta_{h,T})}{\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}}\leq\frac{z_{1-a/2}}{T^{1/2-\zeta}}\right)$
    $=P\left(0\leq\frac{\operatorname{aBias}(\hat{\delta}_{h})}{\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}}\leq 0\right)=0,$
    where the last equality uses the fact that $\operatorname{aBias}(\hat{\delta}_{h})\neq 0$.
  2. 2.
    $\zeta>1/2$: In this case
[…]
    and thus
    $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-a.$
  3. 3.
    $\zeta=1/2$: In this case
[…]
    Thus,
    $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\beta}_{h}))=1-r(b_{h},z_{1-a/2}),$
    where $b_{h}\equiv\operatorname{aBias}(\hat{\delta}_{h})/\sqrt{\operatorname{aVar}(\hat{\delta}_{h})}$ and $r(b;c)\equiv P_{Z\sim N(0,1)}(|Z+b|>c)=\Phi(-c-b)+\Phi(-c+b)$.
In summary, $\lim_{T\to\infty}P(\theta_{h,T}\in\operatorname{CI}(\hat{\delta}_{h}))=\lim_{T\to\infty}\{1-r(T^{1/2-\zeta}b_{h};z_{1-a/2})\}$. ∎
C.3 Proof of 3.3
@@
It now follows as in the proof of 3.1 that $\operatorname{aVar}(\hat{\beta}_{h})=\operatorname{aVar}(\hat{\delta}_{h})$. Then Proposition 4.1 implies that $\operatorname{aBias}(\hat{\delta}_{h})=0$. ∎
C.4 Proof of 4.1
@@
C.5 Proof of 4.2
Write $\hat{\theta}_{h}(\omega)=\hat{\delta}_{h}+\omega(\hat{\beta}_{h}-\hat{\delta}_{h})$. By Proposition 3.3, the two terms are asymptotically independent of each other, and the second term has asymptotic variance $\omega^{2}\{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})\}$. Hence,

### v2

@@ Appendix B Proofs
$\sum_{\ell=1}^{\infty}\Upsilon_{\ell}\Upsilon_{\ell}^{\prime}=\operatorname{aVar}(\hat{\boldsymbol{\beta}})-\operatorname{aVar}(\hat{\boldsymbol{\delta}}).$
Proof.
By definition of $\Upsilon_{\ell}$, it suffices to show that, for any indices $a,b\in\{1,\dots,k\}$,
$\sum_{\ell=1}^{\infty}\operatorname*{vec}(\Xi_{i_{a}^{*},j_{a}^{*},h_{a},\ell})^{\prime}\operatorname*{vec}(\Xi_{i_{b}^{*},j_{b}^{*},h_{b},\ell})=\operatorname{aCov}(\hat{\beta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\beta}_{i_{b}^{*},j_{b}^{*},h_{b}})-\operatorname{aCov}(\hat{\delta}_{i_{a}^{*},j_{a}^{*},h_{a}},\hat{\delta}_{i_{b}^{*},j_{b}^{*},h_{b}}).$ (B.2)
Multiplying out terms, we find that the left-hand side above equals
$\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{i_{a}^{*},j_{a}^{*},h_{a},\ell}^{\prime}\Xi_{i_{b}^{*},j_{b}^{*},h_{b},\ell})$ $=\sum_{\ell=1}^{\infty}\operatorname*{trace}\left(A^{\ell-1}\Sigma(A^{\prime})^{\ell-1}S^{-1}\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right)$
@@
$\qquad-\sum_{\ell=1}^{h_{b}}\operatorname*{trace}\left(A^{\ell-1}H_{\bullet,j_{b}^{*}}e_{i_{b}^{*},n}^{\prime}A^{h_{b}-\ell}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right)$
@@
$\qquad-\sum_{\ell=1}^{h_{a}}\operatorname*{trace}\left(A^{\ell-1}H_{\bullet,j_{a}^{*}}e_{i_{a}^{*},n}^{\prime}A^{h_{a}-\ell}\Sigma\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}^{\prime}S^{-1}\right)$
$\qquad+\mathbbm{1}(j_{a}^{*}=j_{b}^{*})\sigma_{j_{a}^{*}}^{-2}\sum_{\ell=1}^{\min\{h_{a},h_{b}\}}e_{i_{b}^{*},n}^{\prime}A^{h_{b}-\ell}\Sigma(A^{\prime})^{h_{a}-\ell}e_{i_{a}^{*},n}.$
We now evaluate each of the four terms on the right-hand side above. The first term equals
$\operatorname*{trace}\Bigg{(}\underbrace{\sum_{\ell=1}^{\infty}A^{\ell-1}\Sigma(A^{\prime})^{\ell-1}}_{=S}S^{-1}\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\Bigg{)}=\operatorname*{trace}\left(\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right).$
The second term (in the earlier display) equals
$-\operatorname*{trace}\Bigg{(}\underbrace{\sum_{\ell=1}^{h_{b}}A^{\ell-1}H_{\bullet,j_{b}^{*}}e_{i_{b}^{*},n}^{\prime}A^{h_{b}-\ell}}_{=\sum_{\ell=1}^{h_{b}}A^{h_{b}-\ell}H_{\bullet,j_{b}^{*}}e_{i_{b}^{*},n}^{\prime}A^{\ell-1}=\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\Bigg{)}=-\operatorname*{trace}\left(\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right),$
and the third term (in the earlier display) also equals this quantity by a symmetric calculation. In conclusion, we have shown
$\sum_{\ell=1}^{\infty}\operatorname*{trace}(\Xi_{i_{a}^{*},j_{a}^{*},h_{a},\ell}^{\prime}\Xi_{i_{b}^{*},j_{b}^{*},h_{b},\ell})$
$=\mathbbm{1}(j_{a}^{*}=j_{b}^{*})\sigma_{j_{a}^{*}}^{-2}\sum_{\ell=1}^{\min\{h_{a},h_{b}\}}e_{i_{b}^{*},n}^{\prime}A^{h_{b}-\ell}\Sigma(A^{\prime})^{h_{a}-\ell}e_{i_{a}^{*},n}-\operatorname*{trace}\left(\Psi_{i_{b}^{*},j_{b}^{*},h_{b}}\Sigma\Psi_{i_{a}^{*},j_{a}^{*},h_{a}}^{\prime}S^{-1}\right).$
The desired result (B.2) now follows from Corollary A.2. ∎
B.4 Proof of Corollary 3.2
@@
It now follows as in the proof of Corollary A.2 that $\operatorname{aVar}(\hat{\beta}_{h})=\operatorname{aVar}(\hat{\delta}_{h})$. Then Proposition 4.1 implies that $\operatorname{aBias}(\hat{\delta}_{h})=0$. ∎
B.5 Proof of Corollary 4.1
@@
B.6 Proof of Corollary 4.2
Write $\hat{\theta}_{h}(\omega)=\hat{\delta}_{h}+\omega(\hat{\beta}_{h}-\hat{\delta}_{h})$. By Corollary A.2, the two terms are asymptotically independent of each other, and the second term has asymptotic variance $\omega^{2}\{\operatorname{aVar}(\hat{\beta}_{h})-\operatorname{aVar}(\hat{\delta}_{h})\}$. Hence,